# Efficient sequential Bayesian inference for state-space epidemic models using ensemble data assimilation

### v1

@@ 3.2 Application to the 2022 Monkeypox outbreak in the USA
We applied the proposed methodology to infer the transmission dynamics of the 2022 U.S. monkeypox (mpox) outbreak, caused by the West African Clade II of the monkeypox virus. This outbreak represented an unprecedented global event, with the United States reporting among the highest case counts worldwide. The typical incubation period ranges from 3213-21 days, and infectiousness generally lasts 242-4 weeks. The dataset consists of daily confirmed mpox case counts reported by the U.S. Centers for Disease Control and Prevention (CDC) from 6 May to 31 December 2022. Recent studies have applied Bayesian filtering to mpox data. For instance, Saldaña et al. (2023) estimated the instantaneous reproduction number and growth rates using sequential Bayesian updating, while Papageorgiou & Kolias (2024) proposed a particle filter incorporating penalty factors specific to mpox transmission. Our framework offers a complementary perspective by enabling scalable joint inference of latent epidemic states and model parameters, allowing uncertainty quantification directly from noisy daily surveillance data.
Given the pronounced day-to-day variability in the raw counts, we employed the Negative Binomial observation model as the primary specification. Preliminary experiments with an equidispersed (Poisson) model led to unstable posterior trajectories and parameter degeneracy (results not shown). The Poisson assumption becomes more plausible when incidence was aggregated over 77-day intervals, as temporal averaging reduces excess variability (see Figure F.1 in the Appendix F).
@@
Figure 7 shows the estimated daily incidence and $R_{\text{eff}}(t)$ over time. The filtered mean trajectories closely track the observed case counts, while the credible bands reflect the uncertainty associated with both stochastic epidemic dynamics and reporting noise. The estimate of $R_{\text{eff}}(t)$ indicates values well above one during the early growth phase, followed by a steady decline after the declaration of the state of emergency, and eventually stabilizing below one as the epidemic subsides. The estimated latent SEIR trajectories and transmission rates are shown in Figure F.2 in the Appendix F.
Figure 8 presents the posterior trajectories of key epidemiological parameters, with corresponding priors and posterior summaries reported in Table 2. As more data are assimilated, these parameters become increasingly well-identified, as reflected by the progressive narrowing of the posterior intervals. The analysis indicates a mean incubation period of approximately $5.3$ days (95% CrI: 4.64.66.26.2 days), consistent with early empirical estimates reported by Madewell et al. (2023), and a mean infectious period of approximately $18.2$ days (95% CrI: 15.215.222.222.2 days). The overdispersion parameter $\phi$ exhibits posterior mass distinctly separated from zero, confirming substantial extra-Poisson variability. For comparison, we also implemented the standard SMC2 algorithm under identical model and prior settings. Results in Figure F.3 in the Appendix F show that both methods produce a very similar evolution of the filtering distributions for the parameters, indicating that the ensemble-based approximation in eSMC2 accurately captures posterior uncertainty. Notably, A single run with these settings required approximately 16 minutes of CPU time, compared with nearly 2.5 hours for the equivalent SMC2 implementation, demonstrating the efficiency gains achieved by the EnKF-based approach.
@@
$\beta_{0}$ Initial condition transmission rate $\mathcal{U}(0.2,0.3)$ -
@@
Figure 9: Posterior-predictive forecasts at different starting dates. Filtering estimates are shown in blue, and forecast distributions in purple, with corresponding 50%, 75%, 90%, and 95% credible intervals. In-sample observations (prior to the forecast start date) are shown in red, and out-of-sample observations in yellow. Vertical orange dashed lines indicate the forecast start dates.
Figure 9 presents posterior-predictive forecasts of the mpox epidemic, initiated at several points during the outbreak and projected over a 14-day horizon under the assumption that the transmission rate $\beta_{t}$ remains fixed at its last estimated value. These forecasts are obtained by propagating the latent state $x_{T}$ forward through the state-space transition dynamics and then sampling synthetic observations $\tilde{y}_{T+k}$ from the corresponding observation model, conditional on the projected states. In most cases, the forecast distributions captured the subsequent decline in incidence, with credible intervals widening appropriately as the horizon increased. However, the earliest forecast (top-left panel) failed to anticipate the turning point and instead projected continued growth, reflecting the limitation of assuming a constant transmission rate during periods of rapid change. We retain this simplifying assumption, as propagating the Gaussian-process dynamics of $\beta_{t}$ forward, without observations to perform the update stage, could produce unstable or unrealistic trajectories.
Figure 10: Long-term forecast performance. Again, forecast distribution with credible intervals for the final phase of the outbreak is in purple. In-sample data (red) are shown before the forecast start date (vertical dashed line), while out-of-sample data (yellow) are used for validation.
Figure 10 focuses on long-term forecasts during the final epidemic phase. The model successfully reproduced the overall downward trend, though fluctuations in daily reports were harder to capture. Forecast evaluation indicates good calibration with the 95% posterior credible interval encompassing all out-of-sample observations, and the RMSE was 4.179. These results demonstrate that eSMC2 can provide calibrated and accurate probabilistic forecasts, offering valuable support for real-time epidemic monitoring and decision-making.
@@ 4 Discussion and limitations
In this work, we investigated sequential Bayesian inference for latent states and parameters in compartmental models observed through incidence data. We proposed the eSMC2, an extension of the SMC2 framework that replaces the particle filter with an EnKF-derived likelihood, enabling computationally efficient sequential updates of both latent epidemic trajectories and model parameters. This approach builds on the eMCMC method of Drovandi et al. (2022) but extends it to a fully sequential setting, allowing time-resolved inference in dynamic epidemic systems. By leveraging the ensemble approximation of the likelihood, eSMC2 reduces the computational burden while retaining the ability to capture nonlinear state evolution.
@@
Methodologically, the EnKF provides a scalable alternative to standard particle filtering, particularly in moderately high-dimensional or computationally constrained settings. While the EnKF likelihood is only a Gaussian approximation, we employ an unbiased correction, following Drovandi et al. (2022), which reduces sensitivity to the ensemble size. In addition, rather than assuming a fixed observation noise variance, we dynamically update the observation noise variance at each filtering step using the ensemble-predicted state estimates. This adaptive treatment of observation uncertainty allows the filter to better reflect time-varying variability in incidence data. Although we do not provide formal bias quantification, previous studies indicate that EnKF-based likelihoods yield stable and accurate inference in moderately nonlinear systems when combined with MCMC (Khalil et al. 2015, Drovandi et al. 2022). Our simulation studies confirm this behavior, showing that the approximation error remains small and posterior means closely match those obtained with standard SMC2. Importantly, the diffusion-driven SEIR formulation combined with Bayesian filtering provides a flexible mechanism for detecting temporal changes in transmission, such as those resulting from behavioral shifts or intervention measures, without requiring explicit change-point modeling. In our simulations, the method accurately recovers both the transmission rate and the effective reproduction number, even without knowledge of their true trajectories. The application to the 2022 U.S. mpox outbreak further demonstrates that eSMC2 delivers well-calibrated, near-real-time estimates of transmission dynamics and effective reproduction numbers, supporting timely situational awareness and public health decision-making.
@@
Despite its practical and computational advantages, several methodological and theoretical limitations should be noted. Unlike standard SMC2, which is theoretically exact in the limit of infinite state and parameter particles, eSMC2 inherits bias from the EnKF likelihood approximation. This bias may become more pronounced in strongly nonlinear or multimodal epidemic systems, and rigorous convergence guarantees for nonlinear, non-Gaussian settings remain an open challenge. While convergence of the EnKF likelihood has been established in linear-Gaussian systems with rate $\mathcal{O}(N_{x}^{-1/2})$ (Chen et al. 2022), extending such results to nonlinear systems is nontrivial. Although the proposed method is sequential, it is not fully “online”, as each call of the PMMH kernel still involves running the EnKF on all data available up to that point. One potential avenue to further improve computational efficiency would be to develop a fully recursive implementation analogous to a Nested Particle Filter (NPF; Crisan & Míguez 2018), in which the inner particle filter could be replaced by an EnKF. In our SEIR model, we used only a single data stream (daily incidence). Adding data such as hospitalizations, deaths, or serological surveys could help disentangle the reporting rate from transmission parameters and improve parameter identifiability (Swallow et al. 2022). Finally, while the framework is broadly applicable to other state-space models beyond epidemiology, its performance may vary depending on the degree of nonlinearity, the suitability of the Gaussian approximation for the forecast ensemble and the choice of observation noise. Future research will explore these limitations in more detail.
Data and code availability
Code implementing the eSMC2 algorithm for Bayesian inference in state-space epidemiological models can be downloaded from https://github.com/Dhorasso/eSMC2_Epi_Inference. The 2022 U.S. monkeypox incidence data used in this study are freely accessible from the Centers for Disease Control and Prevention (CDC) at: https://www.cdc.gov/monkeypox/data-research/cases/index.html.
Acknowledgements
@@ A.1 Transition density of the SEIR model
$p(x_{t}|x_{t-1},\theta)=\delta_{f(x_{t-1},\theta)}(S_{t},E_{t},I_{t},R_{t},Z_{t})\,\mathcal{N}(\log(\beta_{t});\penalty 10000\ \log(\beta_{t-1}),\penalty 10000\ \nu_{\beta}^{2}).$ (27)
A.2 Kalman gain and bservation error variance
We consider the stochastic process $(x_{t},y_{t})_{t\geq 0}$ defined in Section 2.1, such that $y_{t}|x_{t}\sim p(y_{t}|x_{t})$ with finite conditional mean $\mathrm{E}[y_{t}|x_{t}]=Hx_{t}$. Then
@@ A.2 Kalman gain and bservation error variance
$\approx\widehat{\Sigma}_{t|t-1}H^{\top}\left(H\widehat{\Sigma}_{t|t-1}H^{\top}+\frac{1}{N_{x}}\sum_{i=1}^{N_{x}}\mathrm{Var}\big[y_{t}|x_{t}^{(f,i)}\big]\right)^{-1}.$ (29)
@@ Appendix C Additional results from simulated experiments
This section presents supplementary figures supporting the results discussed in the main text. Figures C.1C.2 show the filtered estimated trajectories of the unobserved SEIR compartments, while Figures C.3C.4 display the posterior pairwise parameter distributions obtained from SMC2 and eSMC2 for Examples 1 and 2.
@@ Appendix D Sensitivity to ensemble size and prior
We also assessed the impact of using non-informative priors by assigning $\alpha,\gamma,\nu_{\beta}\sim\mathcal{U}(0,1)$ and fitting the model to simulated data. With flat priors, the posterior distribution is largely determined by the likelihood, causing the posterior mode to align closely with the maximum likelihood estimate. Figures D.2 and D.3 summarize the results for Examples 1 and 2, respectively, showing the evolution of the filtered parameter trajectories and the final marginal and joint posterior distributions.
@@ Appendix E Comparison with the Liu and West filter
To further evaluate the performance of eSMC2, we compare it with the well-established particle filtering approach of Liu & West (2001), commonly known as the Liu and West filter. This algorithm provides a benchmark for joint sequential state and parameter estimation within state-space models. It extends the standard particle filter by augmenting the state vector with static parameters and applying kernel density shrinkage to mitigate parameter degeneracy. The algorithm generates samples $\{x_{t}^{(i)},\theta_{t}^{(i)}\}_{i=1}^{N_{x}}\sim p(x_{t},\theta|y_{1:t})$.
@@
The shrinkage parameter $\lambda$ is defined as $\lambda=\sqrt{1-h^{2}}$, where $h^{2}=1-\left(\frac{3\delta-1}{2\delta}\right)^{2}$ is the kernel smoothing parameter. This shrinkage step centers each kernel component around the global mean $\bar{\theta}_{t-1}$ and ensures that the mixture variance matches the posterior variance, thereby preventing over-dispersion of the particle cloud. Following Liu & West (2001), we set $\delta=0.99$ and used $N_{x}=20{,}000$ particles. The prior specifications are identical to those described in Section 3.1.
Figures E.1 and E.2 compare filtered estimates for Examples 1 and 2 obtained with the Liu and West filter and eSMC2. Both methods capture the broad temporal dynamics of $\beta_{t}$, but differences arise in the inferred parameters. In particular, the Liu and West filter consistently overestimates the recovery rate $\gamma$, even with a large ensemble, likely due to the strong correlation between $\gamma$ and the time-varying transmission rate that kernel-based shrinkage cannot fully account for. It also produces wider and less stable credible bands, reflecting increased Monte Carlo variability and weight degeneracy. In contrast, eSMC2 yields posterior means closer to the true values with narrower, more coherent uncertainty intervals, highlighting its improved stability and accuracy in sequential updating.
Figure E.1: Comparison between the Liu and West filter and eSMC2 for simulated data (Example 1). Posterior distributions (filtered means and $95\%$ credible intervals) over time.
Figure E.2: Comparison between the Liu and West filter and eSMC2 for simulated data (Example 2). Posterior distributions (filtered means and $95\%$ credible intervals) over time.

### v2

@@ 3.2 Application to the 2022 Monkeypox outbreak in the USA
We applied the proposed methodology to infer the transmission dynamics of the 2022 U.S. monkeypox (mpox) outbreak, caused by the West African Clade II of the monkeypox virus. This outbreak represented an unprecedented global event, with the United States reporting among the highest case counts worldwide. The typical incubation period ranges from 3213-21 days, and infectiousness generally lasts 242-4 weeks. The dataset consists of daily confirmed mpox case counts reported by the U.S. Centers for Disease Control and Prevention (CDC) from 6 May to 31 December 2022. Recent studies have applied Bayesian filtering to mpox data. For instance, Saldaña et al. (2023) estimated the instantaneous reproduction number and growth rates using sequential Bayesian updating, while Papageorgiou and Kolias (2024) proposed a particle filter incorporating penalty factors specific to mpox transmission. A related stochastic Kalman-based approach, validated on the 2022 Czech Republic mpox data, further discusses the challenges of Bayesian filtering for epidemic modeling with dynamic parameters (Papageorgiou et al., 2025). Our framework offers a complementary perspective by enabling scalable joint inference of latent epidemic states and model parameters, allowing uncertainty quantification directly from noisy daily surveillance data.
Given the pronounced day-to-day variability in the raw counts, we employed the Negative Binomial observation model as the primary specification. Preliminary analyses with an equidispersed (Poisson) observation model produced unstable posterior trajectories and parameter degeneracy, indicating that the Poisson assumption was inconsistent with the variability present in the daily data. The Poisson model becomes more plausible when incidence is aggregated over 7-day intervals, as temporal averaging reduces excess variability. Supporting results for the aggregated analysis are provided in Appendix F.
@@
Fig 9 shows the estimated daily incidence and $R_{\text{eff}}(t)$ over time. The filtered mean trajectories closely track the observed case counts, while the credible bands reflect the uncertainty associated with both stochastic epidemic dynamics and reporting noise. The estimate of $R_{\text{eff}}(t)$ indicates values well above one during the early growth phase, followed by a steady decline after the declaration of the state of emergency, and eventually stabilizing below one as the epidemic subsides. The corresponding latent SEIR trajectories and transmission rates are provided in Appendix F.
@@
Fig 10 presents the posterior trajectories of key epidemiological parameters, with corresponding priors and posterior summaries reported in Table 2. As more data are assimilated, these parameters become increasingly well-identified, as reflected by the progressive narrowing of the posterior intervals. The analysis indicates a mean incubation period of approximately $5.3$ days (95% CrI: 4.64.66.26.2 days), consistent with early empirical estimates reported by Madewell et al. (2023), and a mean infectious period of approximately $18.2$ days (95% CrI: 15.215.222.222.2 days). The overdispersion parameter $\phi$ exhibits posterior mass distinctly separated from zero, confirming substantial extra-Poisson variability. For comparison, we also implemented the standard SMC2 algorithm under identical model and prior settings. Results reported in Appendix F show that both methods produce a very similar evolution of the filtering distributions for the parameters, indicating that the ensemble-based approximation in eSMC2 accurately captures posterior uncertainty. Notably, A single run with these settings required approximately 16 minutes of CPU time, compared with nearly 2.5 hours for the equivalent SMC2 implementation, demonstrating the efficiency gains achieved by the EnKF-based approach.
@@
$\beta_{0}$ Initial condition transmission rate $\mathcal{U}(0.2,\,0.3)$
@@
Fig 11 presents 14-day-ahead posterior-predictive forecasts initiated at multiple time points during the outbreak, assuming the transmission rate $\beta_{t}$ remains fixed at its last estimated value. Forecasts are obtained by propagating the latent state forward through the state-space dynamics and sampling synthetic observations from the observation model. The purpose of this analysis is to evaluate the short-term predictive performance of the proposed eSMC2 framework. We compare the results with an autoregressive (AR) model (Appendix F). The top panel of Fig 11 shows forecasts from eSMC2 and the bottom panel those from the AR model. Quantitative performance is summarised in Table 3, which reports (the) mean absolute error (MAE), weighted interval score (WIS; based on the 50%, 75%, 90%, and 95% intervals), and empirical 95% prediction interval (PI) coverage for each forecast start date. Both models struggle to anticipate the turning point in early August, and at this stage the autoregressive model achieves lower forecast error. However, from mid-September onward, eSMC2 achieves lower MAE and WIS in most forecast windows and maintains coverage close to the nominal 95% level, whereas the autoregressive model exhibits declining coverage in later periods (Table 3). These results indicate that explicitly modelling latent transmission dynamics improves probabilistic calibration and overall forecast accuracy once incidence begins to stabilise or decline. Long-term forecasts for the final epidemic phase are presented in Fig. F4. The model reproduces the overall decline and achieves full 95% predicted interval coverage.
Figure 11: Posterior-predictive forecasts at different starting dates. The top panel shows eSMC2 forecasts, and the bottom panel shows AR forecasts. Median forecasts are displayed in purple (eSMC2) and green (AR), with corresponding 50%, 75%, 90%, and 95% predictive intervals. Observed case counts are shown in red. Vertical orange dashed lines indicate the forecast start dates.
Table 3: Evaluation of the 2 weeks-ahead forecasts of reported mpox cases during the 2022 U.S. outbreak. Bold indicates better-performing model.
eSMC2 AR
Forecast date 95% PI coverage MAE WIS 95% PI coverage MAE WIS
2022-08-04 0.43 230.6 269.5 0.57 204.2 251.6
2022-08-18 1.00 78.5 121.3 1.00 52.1 94.2
2022-09-01 0.86 41.2 74.5 0.93 30.8 60.3
2022-09-15 1.00 25.4 43.9 0.79 41.2 55.5
2022-09-29 1.00 10.7 22.6 0.86 15.2 25.2
2022-10-13 1.00 7.8 15.2 0.79 18.3 22.8
2022-10-27 1.00 5.2 11.2 0.86 7.5 11.7
2022-11-10 1.00 4.6 8.5 0.64 7.3 9.6
2022-11-24 0.93 3.9 6.5 0.57 5.9 7.7
@@ 4 Discussion and limitations
In this work, we investigated sequential Bayesian inference for latent states and parameters in compartmental models observed through incidence data. We proposed the eSMC2, an extension of the SMC2 framework that replaces the particle filter with an EnKF-derived likelihood, enabling computationally efficient sequential updates of both latent epidemic trajectories and model parameters. This approach builds on the eMCMC method of Drovandi et al. (2022) but extends it to a fully sequential setting, allowing time-resolved inference in dynamic epidemic systems. By leveraging the ensemble approximation of the likelihood, eSMC2 reduces the computational burden while retaining the ability to capture nonlinear state evolution. Within the broader landscape of Bayesian calibration approaches, including likelihood-free methods such as ABC (Li et al., 2025), eSMC2 represents a complementary strategy. Likelihood-free approaches are particularly useful when the likelihood is unavailable or prohibitively expensive to evaluate. In contrast, eSMC2 relies on an explicit state-space formulation and an approximated likelihood, enabling principled sequential updating of both latent states and parameters as new data arrive. Consequently, eSMC2 is especially suited to time-series epidemic models where inference is performed sequentially, whereas likelihood-free methods may be preferable for models that do not admit a tractable state-space representation.
@@
Methodologically, the EnKF provides a scalable alternative to standard particle filtering, particularly in moderately high-dimensional or computationally constrained settings. While the EnKF likelihood is only a Gaussian approximation, we employ an unbiased correction, following Drovandi et al. (2022), which reduces sensitivity to the ensemble size. In addition, rather than assuming a fixed observation noise variance, we dynamically update the observation noise variance at each filtering step using the ensemble-predicted state estimates. This adaptive treatment of observation uncertainty allows the filter to better reflect time-varying variability in incidence data. Although we do not provide formal bias quantification, previous studies indicate that EnKF-based likelihoods yield stable and accurate inference in moderately nonlinear systems when combined with MCMC (Khalil et al., 2015; Drovandi et al., 2022). Our simulation studies confirm this behavior, showing that the approximation error remains small and posterior means closely match those obtained with standard SMC2. Importantly, the diffusion-driven SEIR formulation combined with Bayesian filtering provides a flexible mechanism for detecting temporal changes in transmission, such as those resulting from behavioral shifts or intervention measures, without requiring explicit change-point modeling. In our simulations, the method accurately recovers both the transmission rate and the effective reproduction number, even without knowledge of their true trajectories. The application to the 2022 U.S. mpox outbreak further demonstrates that eSMC2 delivers well-calibrated, near-real-time estimates of transmission dynamics and effective reproduction numbers, supporting timely situational awareness and public health decision-making. It should be noted that this orecasting analysis is retrospective and does not account for real-time complexities such as data revisions and reporting delays, so predictive performance may be optimistic relative to an operational setting (Charniga et al., 2024).
@@
Despite its practical and computational advantages, several methodological and theoretical limitations should be noted. Unlike standard SMC2, which is theoretically exact in the limit of infinite state and parameter particles, eSMC2 inherits bias from the EnKF likelihood approximation. This bias may become more pronounced in strongly nonlinear or multimodal epidemic systems, and rigorous convergence guarantees for nonlinear, non-Gaussian settings remain an open challenge. While convergence of the EnKF likelihood has been established in linear-Gaussian systems with rate $\mathcal{O}(N_{x}^{-1/2})$ (Chen et al., 2022), extending such results to nonlinear systems is nontrivial. Although the proposed method is sequential, it is not fully “online”, as each call of the PMMH kernel still involves running the EnKF on all data available up to that point. One potential avenue to further improve computational efficiency would be to develop a fully recursive implementation analogous to a Nested Particle Filter (NPF) (Crisan and Míguez, 2018), in which the inner particle filter could be replaced by an EnKF. From a modeling standpoint, our SEIR application relies on a single data stream (daily incidence). Incorporating additional observations—such as hospitalizations, deaths, or serological survey data could improve parameter identifiability, particularly in disentangling reporting rates from transmission parameters (Swallow et al., 2022). More generally, epidemiological models often include parameters constrained to the unit interval (e.g., reporting or mortality ratios). In such cases, one may either assign a Beta prior, which naturally enforces support on $(0,1)$, or apply a logit transformation and conduct inference on the unconstrained scale. Finally, while the framework is broadly applicable to other state-space models beyond epidemiology, its performance may vary depending on the degree of nonlinearity, the suitability of the Gaussian approximation for the forecast ensemble and the choice of observation noise. Future research will explore these limitations in more detail.
Acknowledgments
@@
Data availability
All code used to produce the analyses in this paper is available at https://github.com/Dhorasso/bma-smc2-dthp-seir. The datasets used are publicly available from the HPSC at https://respiratorydisease-hpscireland.hub.arcgis.com/pages/influenza for influenza data, and from the Irish COVID-19 Data Hub at https://COVID19ireland-geohive.hub.arcgis.com/ for COVID-19 data.
@@ A.1 Transition density of the SEIR model
$p(x_{t}|x_{t-1},\theta)=\delta_{f(x_{t-1},\theta)}(S_{t},E_{t},I_{t},R_{t},Z_{t})\,\mathcal{N}(\log(\beta_{t});~\log(\beta_{t-1}),~\nu_{\beta}^{2}).$ (27)
A.2 Kalman gain and observation error variance
We consider the stochastic process $(x_{t},y_{t})_{t\geq 0}$ defined in Section 2 of the main text, such that $y_{t}|x_{t}\sim p(y_{t}|x_{t})$ with finite conditional mean $\mathrm{E}[y_{t}|x_{t}]=Hx_{t}$. Then
@@ A.2 Kalman gain and observation error variance
$\approx\widehat{\Sigma}_{t|t-1}H^{\top}\left(H\widehat{\Sigma}_{t|t-1}H^{\top}+\underbrace{\frac{1}{N_{x}}\sum_{i=1}^{N_{x}}\mathrm{Var}\big[y_{t}|x_{t}^{(f,i)}}_{V_{t}^{N_{x}}}\big]\right)^{-1}.$ (29)
@@ Appendix C Additional results from simulated experiments
This section presents supplementary figures supporting the results discussed in the main text. Figures C.1C.3 show the filtered estimated trajectories of the unobserved SEIR compartments, while Figures C.4C.6 display the posterior pairwise parameter distributions obtained from SMC2 and eSMC2 for Examples 1, 2 and 3.
@@ Appendix D Sensitivity to ensemble size and prior
We also assessed the impact of using non-informative priors by assigning $\alpha,\nu_{\beta}\sim\mathcal{U}(0,1)$ and $\gamma\sim\mathcal{U}(0,1)$ for Examples 1 and 2, $\gamma\sim\mathcal{U}(0.1,0.2)$ for Example 3, and fitting the model to simulated data. With flat priors, the posterior distribution is largely determined by the likelihood, causing the posterior mode to align closely with the maximum likelihood estimate. Figures D.2D.3 and D.4 summarize the results for Examples 1, 2 and 3, respectively, showing the evolution of the filtered parameter trajectories and the final marginal and joint posterior distributions.
@@ Appendix E Comparison with the Liu and West filter
To further evaluate the performance of eSMC2, we compare it with the well-established particle filtering approach of Liu and West (2001), commonly known as the Liu and West filter. This algorithm provides a benchmark for joint sequential state and parameter estimation within state-space models. It extends the standard particle filter by augmenting the state vector with static parameters and applying kernel density shrinkage to mitigate parameter degeneracy. The algorithm generates samples $\{x_{t}^{(i)},\theta_{t}^{(i)}\}_{i=1}^{N_{x}}\sim p(x_{t},\theta|y_{1:t})$.
@@
The shrinkage parameter $\lambda$ is defined as $\lambda=\sqrt{1-h^{2}}$, where $h^{2}=1-\left(\frac{3\delta-1}{2\delta}\right)^{2}$ is the kernel smoothing parameter. This shrinkage step centers each kernel component around the global mean $\bar{\theta}_{t-1}$ and ensures that the mixture variance matches the posterior variance, thereby preventing over-dispersion of the particle cloud. Following Liu and West (2001), we set $\delta=0.99$ and used $N_{x}=20{,}000$ particles. The prior specifications are identical to those described in Section 3 of the main text.
Figure E.1 compares filtered estimates obtained with the Liu and West filter and eSMC2, along with the final posterior distributions from a long PMCMC run ($10^{4}$ iterations), displayed as boxplots at $T=60$. Both filtering methods track the parameter trajectories, but differences arise in accuracy and uncertainty quantification. The Liu and West filter consistently overestimates the recovery rate $\gamma$, likely due to kernel-based shrinkage introducing bias in the joint estimation of static parameters. It also produces wider and less stable credible bands, reflecting increased Monte Carlo variability and weight degeneracy. In contrast, eSMC2 yields posterior means closer to the true values with narrower, more coherent uncertainty intervals, highlighting its improved stability and accuracy in sequential updating. The PMCMC boxplots serve as a gold-standard reference: eSMC2 credible intervals align closely with the PMCMC posterior at $T=60$, whereas the Liu and West filter shows more pronounced discrepancies, particularly for $\gamma$.
Figure E.1: Comparison between the Liu and West filter, eSMC2, and PMCMC for simulated data in Example 1. Filtered means and $95\%$ credible intervals over time for parameters $\alpha$, $\gamma$, and $\nu_{\beta}$. Boxplots at $T=60$ display the marginal posterior distributions obtained from the PMCMC at time 60.